### v1

@@ Abstract
The $T_{\rm eff}$ range (6200 – 6000 K) of the NGC 188 subgiant progenitors coincides, in part, with the Li–$T_{\rm eff}$ Plateau observed in metal-poor dwarfs [40, 41]. Rotational mixing depleted the Li of these progenitors by about a factor of 10, as suggested by current turnoff stars (Figures 2, 3) and a presumed initial A(Li) near the meteoritic A(Li), or slightly higher to account for Galactic Li production [42, 43] and NGC 188’s super-solar metallicity. It is then plausible that metal-poor dwarfs are depleted by only about a factor of 3, due to their lower mass and potentially lower initial angular momentum. Such a depletion could reconcile the observed level of the Li-$T_{\rm eff}$ plateau with the A(Li) predicted in standard models of Big Bang Nucleosynthesis [44, 45].
Figure 2: A(Li) versus $T_{\rm eff}$ for turnoff stars and subgiants in M67 and NGC 188. The single members (SM), binary members (BM), single likely members (SM?), and binary likely members (BM?) of NGC 188 are shown. Members of M67 from [35], with membership probability $>$ 0.5 from [46], are also shown.
@@ Methods
To focus on the topic of this study, we have carefully selected 96 turnoff and subgiant members from NGC 188, including only turnoff stars more evolved than $V=$15.4 mag, and subgiants as bright as $V=$ 14.16 mag. Assuming a distance modulus (m-M) = 11.48 mag and E(B-V) = 0.09 mag [38], and applying the extinction relation $A_{V}$ = 3.1 E(B-V) [47], we transform the apparent $V$ to $M_{V}$, and show the selected stars in Extended Figure 1. For comparison, we also show members of M67 [35], transforming them assuming (m-M) = 9.75 mag and E(B-V) = 0.04 mag [39]. A study of subgiants and giants brighter than $V=$ 14.16 mag and the apparent production of Li in some of these stars was presented in [38], and main sequence stars fainter than $V=$15.4 mag will be discussed in future work.
@@
The MOOG stellar line analysis program can be accessed from the website: https://www.as.utexas.edu/~chris/moog.html. Stellar model atmospheres are obtained from the following source: http://kurucz.harvard.edu/grids.html.
@@ Acknowledgments
5951 132.88859 11.81431 12.767 0.577 6112 $<$1.39
Notes 1: ID, $V$, and $B-V$ values are sourced from [51]. The original RA and DEC, initially in B1950, have been converted to J2000 using the astropy package [52].
Notes 2: The original $T_{\rm eff}$ and A(Li) values have been recalibrated to our $T_{\rm eff}$ scale, enabling a direct comparison with our dataset. Only stars with M67 cluster membership $>$ 0.5 from [46] are shown.
Extended Figure 1: Hertzsprung-Russell diagram of stars from M67 and NGC 188. Small dots indicate members of M67 and NGC 188, selected based on a membership probability greater than 0.5 from [46]. The $V$ and $B-V$ values for Gaia members in M67 are retrieved by cross-matching with data from [51], and for NGC 188 are by cross-matching with data from [53]. Larger symbols indicate stars selected in this study. Disks show stars with Li detections whereas inverted triangles show stars with Li upper limits. Single members (SM) and single likely members (SM?) of NGC 188 are in orange whereas binary members (BM) and binary likely members (BM?) are in yellow. Members of M67 from [35], with membership probability $>$ 0.5 from [46], are shown in green.

### v2

@@ Abstract
[…]
Figure 2: A(Li) versus $T_{\rm eff}$ for turnoff stars and subgiants in M67 and NGC 188. The single members (SM), binary members (BM), single likely members (SM?), and binary likely members (BM?) of NGC 188 are shown. Members of M67 from [35], with membership probability >> 0.5 from [46], are also shown.
@@ Methods
To focus on the topic of this study, we have carefully selected 96 turnoff and subgiant members from NGC 188, including only turnoff stars more evolved than V=V=15.4 mag, and subgiants as bright as $V=$ 14.16 mag. Assuming a distance modulus (m-M) = 11.48 mag and E(B-V) = 0.09 mag [38], and applying the extinction relation $A_{V}$ = 3.1 E(B-V) [47], we transform the apparent $V$ to $M_{V}$, and show the selected stars in Extended Figure 1. For comparison, we also show members of M67 [35], transforming them assuming (m-M) = 9.75 mag and E(B-V) = 0.04 mag [39]. A study of subgiants and giants brighter than $V=$ 14.16 mag and the apparent production of Li in some of these stars was presented in [38], and main sequence stars fainter than V=V=15.4 mag will be discussed in future work.
@@
The MOOG stellar line analysis program can be accessed from the website: https://www.as.utexas.edu/˜chris/moog.html. Stellar model atmospheres are obtained from the following source: http://kurucz.harvard.edu/grids.html.
@@ Acknowledgments
5951 132.88859 11.81431 12.767 0.577 6112 <<1.39
Notes 1: ID, $V$, and BVB-V values are sourced from [51]. The original RA and DEC, initially in B1950, have been converted to J2000 using the astropy package [52].
Notes 2: The original $T_{\rm eff}$ and A(Li) values have been recalibrated to our $T_{\rm eff}$ scale, enabling a direct comparison with our dataset. Only stars with M67 cluster membership >> 0.5 from [46] are shown.
Extended Figure 1: Hertzsprung-Russell diagram of stars from M67 and NGC 188. Small dots indicate members of M67 and NGC 188, selected based on a membership probability greater than 0.5 from [46]. The $V$ and BVB-V values for Gaia members in M67 are retrieved by cross-matching with data from [51], and for NGC 188 are by cross-matching with data from [53]. Larger symbols indicate stars selected in this study. Disks show stars with Li detections whereas inverted triangles show stars with Li upper limits. Single members (SM) and single likely members (SM?) of NGC 188 are in orange whereas binary members (BM) and binary likely members (BM?) are in yellow. Members of M67 from [35], with membership probability >> 0.5 from [46], are shown in green.